OpenStudy (anonymous):

\[\int\limits_{}^{} \sqrt{t^{2}-1} dt\] u-sub?

OpenStudy (anonymous):

i dont think this is a u sub.... try doing integration by parts

OpenStudy (anonymous):

wait no i dont think that will workx

OpenStudy (anonymous):

I think you want a trig substitution.

OpenStudy (anonymous):

so like maybe the reciprocal of arcsec then?

OpenStudy (anonymous):

cause arcsec is 1/(u^2 - a^2)^.5

OpenStudy (anonymous):

soo maybe if you flip it around for both sides..... you would get the integral of ure EQ.

OpenStudy (anonymous):

just an idea

OpenStudy (anonymous):

\[\int\sqrt{t^2-1}dt = \int sec(t) dt \]

OpenStudy (anonymous):

Beyond that you'll either need to use a table or derive the integral of the sec function.

OpenStudy (anonymous):

just an idea

OpenStudy (anonymous):

Wait, no that's wrong.

OpenStudy (anonymous):

WHYYY